# Random expansions of trees with bounded height

### v1

@@ 7.4. Balance
In this section we prove statement (C), in the first part of Lemma 7.17, but we also prove other results that will be used to prove statement (A).
@@
Suppose that $p_{\rho}(\bar{x},y)$ is a complete self-contained closure type over $\sigma_{\rho}$ and that $p(\bar{x},y)$ is a self-contained closure type over $\sigma$ such that $p\models p_{\rho}$ (or equivalently, $p_{\rho}$ is equivalent to $p\negthickspace\upharpoonright\negthickspace\sigma_{\rho}$). Suppose that $\mathrm{rank}_{y}(p_{\rho})\ (=\mathrm{rank}_{y}(p))=1$ and that for every $R\in\sigma\setminus\sigma_{\rho}$, if $\bar{z}$ is a subsequence of $\bar{x}\bar{y}$ and $p\models R(\bar{z})$ or $p\models\neg R(\bar{z})$, then $\bar{z}$ contains $y$. Furthermore, suppose that, for some $n$, $\bar{a}\in(T_{n})^{|\bar{x}|}$, $\mathcal{A}^{\prime}\in\mathbf{W}_{n}^{\rho}$, $P\subseteq p_{\rho}(\bar{a},\mathcal{A}^{\prime})$, and $\mathbb{P}_{n}^{\rho}(\mathcal{A}^{\prime})>0$. There is $\alpha\in[0,1]$, depending only on $p_{\rho}$ and $p$, such that for every $\varepsilon>0$ there is $c>0$, depending only on $\alpha$ and $\varepsilon$, such that if $n$ and $|P|$ are large enough, then
@@
Proof. We adopt all assumptions of the lemma. Since $\mathbb{P}_{n}^{\rho}(\mathcal{A}^{\prime})>0$ it follows from Lemma 6.10 that $\mathbb{P}_{n}\big{(}\mathbf{W}^{\mathcal{A}^{\prime}}\big{)}>0$. As $\mathcal{A}^{\prime}\models p_{\rho}(\bar{a},b)$ for all $b\in P$, it follows that for all $b\in P$, $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{p_{\rho}(\bar{a},b)}\big{)}>0$. Lemma 7.11 tells that, for some $\alpha$, $(p,p_{\rho})$ converges to $\alpha$ with respect to $\mathbb{G}$. So for every $\varepsilon>0$, if $n$ is large enough we have
@@
By part (i) of Lemma 7.17 we have proved statement (C) above (at the end of Section 7.2).
Although we have completed the inductive step for part (3) of Assumption 7.8 we will continue to prove results about balanced triples because we need them for proving statement (A) and the main results of this article.
@@
Then $\bar{y}=z\bar{u}$ so $s(\bar{x},z,\bar{u})$ is the same as $p(\bar{x},z,\bar{u})$ (and $p_{\tau}$ is the same as $s_{\tau}$). Since $p_{\tau}(\bar{x},z,\bar{u})$ is self-contained it follows that if $\mathcal{T}_{n}\models p_{\tau}(\bar{a},b,\bar{c})$ then $\mathcal{T}^{\prime}:=\mathcal{T}_{n}\negthickspace\upharpoonright\negthickspace\mathrm{rng}(b\bar{c})$ is a subtree of $\mathcal{T}_{n}$ which is rooted in $b$ and the isomorphism type of $\mathcal{T}^{\prime}$ is determined by $p_{\tau}$ alone. It is easy to see that if $\mathcal{T}_{n}\models r_{\tau}(\bar{a},b)$ and $\mathcal{T}^{\prime\prime}$ is a subtree of $\mathcal{T}_{n}$ which is rooted in $b$ and isomorphic to $\mathcal{T}^{\prime}$, then the elements of $T^{\prime\prime}\setminus\{b\}$ can be ordered in a unique way as $\bar{c}$ so that $\mathcal{T}_{n}\models p_{\tau}(\bar{a},b,\bar{c})$. Therefore there is a bijection between $p_{\tau}(\bar{a},b,\mathcal{T}_{n})$ and the subtrees of $\mathcal{T}_{n}$ that are rooted in $b$ and isomorphic to $\mathcal{T}^{\prime}$. Recall that $\mathsf{N}_{n}(b,\mathcal{T}^{\prime}):=\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})$ is the number of subtrees of $\mathcal{T}_{n}$ that are rooted in $b$ and isomorphic to $\mathcal{T}^{\prime}$. Now we have $|p_{\tau}(\bar{a},b,\mathcal{T}_{n})|=\mathsf{N}_{n}(b,\mathcal{T}^{\prime})$ if $p_{\tau}(\bar{a},b,\mathcal{T}_{n})\neq\emptyset$. Recall that, by Remark 6.4, we have $\mathsf{N}_{n}(b,\mathcal{T}^{\prime})\geq g_{1}(n)$ whenever $n$ is large enough and $\mathsf{N}_{n}(b,\mathcal{T}^{\prime})>0$.
@@
We now assume that $p_{\tau}(\bar{a},\mathcal{T}_{n})\neq\emptyset$; otherwise there is nothing to prove. Then $r_{\tau}(\bar{a},\mathcal{T}_{n})\neq\emptyset$ and for every $b\in r_{\tau}(\bar{a},\mathcal{A})$ we have $|p_{\tau}(\bar{a},b,\mathcal{T}_{n})|=\mathsf{N}_{n}(b,\mathcal{T}^{\prime})\geq g_{1}(n)$ if $n$ is large enough. We will now define a partition $P_{1},\ldots,P_{\mu}$ of $r_{\tau}(\bar{a},\mathcal{T}_{n})$. Recall the functions $g_{1},g_{2},g_{3}:\mathbb{N}\to\mathbb{R}^{+}$ from Assumption 6.3.
@@
    Let $\lambda\geq 1$ be the minimal integer such that there is $b\in r_{\tau}(\bar{a},\mathcal{A})$ such that $g_{1}(n)+(\lambda-1)g_{2}(n)\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<g_{1}(n)+\lambda g_{2}(n)$.
@@
      If $\lambda=1$ then let $P_{1}$ contain all $b\in r_{\tau}(\bar{a},\mathcal{T}_{n})$ such that $m_{1,1}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{1,2}$ where $m_{1,1}:=g_{1}(n)$ and $m_{1,2}:=m_{1,1}+(\lambda+3)g_{2}(n)$.
@@
      If $\lambda>1$ then let $P_{1}$ contain all $b\in r_{\tau}(\bar{a},\mathcal{T}_{n})$ such that $m_{1,1}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{1,2}$ where $m_{1,1}:=g_{1}(n)+(\lambda-1)g_{2}(n)$ and $m_{1,2}:=m_{1,1}+(\lambda+3)g_{2}(n)$.
@@
    Now suppose that the parts $P_{1},\ldots,P_{l}$ have been defined such that $|P_{i}|\geq g_{3}(n)$ for all $i$ and that numbers $m_{i,j}$ (with $j\in\{1,2\}$) have been defined such that $P_{i}$ contains all $b\in r_{\tau}(\bar{a},\mathcal{A})$ such that $m_{i,1}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{i,2}$, and $m_{i+1,1}=m_{i,2}$ if $i<l$.
    Let $\lambda\geq 1$ be the minimal integer (if it exists) such that there is $b\in r_{\tau}(\bar{a},\mathcal{A})$ such that $m_{l,2}+\lambda g_{2}(n)\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{l,2}+(\lambda+1)g_{2}(n)$.
@@
      If $\lambda=1$ then let $P_{l+1}$ contain all $b\in r_{\tau}(\bar{a},\mathcal{A})$ such that $m_{l+1,1}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{l+1,2}$ where $m_{l+1,1}:=m_{l,2}$ and $m_{l+1,2}:=m_{l,2}+3g_{2}(n)$.
@@
        let $P_{l+1}$ contain all $b\in r_{\tau}(\bar{a},\mathcal{T}_{n})$ such that $m_{l+1,1}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{l+1,2}$ where $m_{l+1,1}:=m_{l,2}+(\lambda-1)g_{2}(n)$ and $m_{l+1,2}:=m_{l,2}+(\lambda+2)g_{2}(n)$, and
@@
        redefine $P_{l}$ so that it contains contain all $b\in r_{\tau}(\bar{a},\mathcal{T}_{n})$ such that $m_{l,1}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{l,2}+g_{2}(n)$ and then redefine $m_{l,2}$ by $m_{l,2}:=m_{l,2}+g_{2}(n)$.
@@
Observe that when a part $P_{i}$ was first defined then, for some real number $m_{i}$, all $b\in P_{i}$ satisfy $m_{i}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{i}+3g_{2}(n)$. Also, each part $P_{i}$ was redefined at most once and in such a way that (with the same $m_{i}$) all $b\in P_{i}$ satisfy $m_{i}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{i}+4g_{2}(n)$. By the construction we have $m_{i}\geq g_{1}(n)$ for all $i$. It follows that
@@
for all $b\in P_{i}$, $m_{i}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})\leq m_{i}+4g_{2}(n)$.
@@
Recall that (by Assumption 6.3) every element of $\mathcal{T}_{n}$ has at most $f_{4}(n)$ children, where $f_{4}$ is a polynomial. Since all members of $r_{\tau}(\bar{a},\mathcal{T}_{n})$ are children of $a_{1}$ it follows that $|r_{\tau}(\bar{a},\mathcal{T}_{n})|\leq f_{4}(n)$. As $P_{1},\ldots,P_{\mu}$ is a partition of $r_{\tau}(\bar{a},\mathcal{T}_{n})$ it follows that $\mu\leq f_{4}(n)$. Let
@@
Proof. Let $\mathcal{A}\in\mathbf{X}_{n}^{\varepsilon}$, so $\mathcal{A}\in\mathbf{X}_{n}^{\varepsilon,i}$ for all $i=1,\ldots,\mu$, and let $\bar{a}\in(T_{n})^{|\bar{x}|}$. Suppose that $r_{\tau}(\bar{a},\mathcal{A})\neq\emptyset$, since otherwise the inequalities are trivial. Fix any $i\in\{1,\ldots,\mu\}$. By (7.9) there is a number $m_{i}\geq g_{1}(n)$ such that if $b\in P_{i}$ then $m_{i}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{i}+4g_{2}(n)$ and hence $m_{i}\leq|p_{\tau}(\bar{a},b,\mathcal{A})|<m_{i}+4g_{2}(n)$.
@@
Let $p(\bar{x},\bar{y})$ and $r(\bar{x},\bar{y})$ be (possibly incomplete) closure types over $\sigma$, where $r$ is $\bar{y}$-positive with respect to $\mathbb{G}$, and let $q(\bar{x})$ be complete closure type over $\sigma$, Then $(p,r,q)$ is balanced with respect to $\mathbb{G}$.
@@
Let $p(\bar{x},\bar{y})$, $r(\bar{x},\bar{y})$, and $q(\bar{x})$ be as in Proposition 7.24 and suppose that $p\wedge r\wedge q$ is cofinally satisfiable. By analysing the proofs of Section 7.4 (this section) we see that if $(p,r,q)$ is 00-balanced, then $(p,r)$ converges to 00. If $(p,r)$ is also eventually constant, then it must be eventually constant with value 00. This implies that for all sufficiently large $n$ and all $\bar{a}\in(T_{n})^{|\bar{x}|}$, $p(\bar{a},\mathcal{A})\cap r(\bar{a},\mathcal{A})=\emptyset$.

### v2

@@ 7.4. Balance
In this section we prove statement (B), in the first part of Lemma 7.17, but we also prove other results that will be used to prove statement (C).
@@
Suppose that $p_{\rho}(\bar{x},y)$ is a complete self-contained closure type over $\sigma_{\rho}$ and that $p(\bar{x},y)$ is a self-contained closure type over $\sigma$ such that $p\models p_{\rho}$ (or equivalently, $p_{\rho}$ is equivalent to $p\negthickspace\upharpoonright\negthickspace\sigma_{\rho}$). Suppose that $\mathrm{rank}_{y}(p_{\rho})\ (=\mathrm{rank}_{y}(p))=1$ and that for every $R\in\sigma\setminus\sigma_{\rho}$, if $\bar{z}$ is a subsequence of $\bar{x}y$ and $p\models R(\bar{z})$ or $p\models\neg R(\bar{z})$, then $\bar{z}$ contains $y$. Furthermore, suppose that $n\in\mathbb{N}^{+}$, $\bar{a}\in(T_{n})^{|\bar{x}|}$, $\mathcal{A}^{\prime}\in\mathbf{W}_{n}^{\rho}$, $P\subseteq p_{\rho}(\bar{a},\mathcal{A}^{\prime})$ and $\mathbb{P}_{n}^{\rho}(\mathcal{A}^{\prime})>0$. There is $\alpha\in[0,1]$, depending only on $p_{\rho}$ and $p$, such that for every $\varepsilon>0$ there is $c>0$, depending only on $\alpha$ and $\varepsilon$, such that if $n$ and $|P|$ are large enough, then
@@
Proof. We adopt all assumptions of the lemma. Since $\mathbb{P}_{n}^{\rho}(\mathcal{A}^{\prime})>0$, it follows from Lemma 6.10 that $\mathbb{P}_{n}\big{(}\mathbf{W}^{\mathcal{A}^{\prime}}\big{)}>0$. As $\mathcal{A}^{\prime}\models p_{\rho}(\bar{a},b)$ for all $b\in P$, it follows that for all $b\in P$, $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{p_{\rho}(\bar{a},b)}\big{)}>0$. Lemma 7.11 tells that, for some $\alpha$, $(p,p_{\rho})$ converges to $\alpha$ with respect to $\mathbb{G}$. So for every $\varepsilon>0$, if $n$ is large enough we have
@@
By part (i) of Lemma 7.17 we have proved statement (B) above (at the end of Section 7.2).
Although we have completed the inductive step for part (3) of Assumption 7.8 we will continue to prove results about balanced triples because we need them for proving statement (C) and the main results of this article.
@@
Then $\bar{y}=z\bar{u}$ so $s(\bar{x},z,\bar{u})$ is equivalent to $p(\bar{x},z,\bar{u})$ (and $p_{\tau}$ is equivalent to $s_{\tau}$). Since $p_{\tau}(\bar{x},z,\bar{u})$ is self-contained it follows that if $\mathcal{T}_{n}\models p_{\tau}(\bar{a},b,\bar{c})$ then $\mathcal{T}^{\prime}:=\mathcal{T}_{n}\negthickspace\upharpoonright\negthickspace\mathrm{rng}(b\bar{c})$ is a subtree of $\mathcal{T}_{n}$ which is rooted in $b$ and the isomorphism type of $\mathcal{T}^{\prime}$ is determined by $p_{\tau}$ alone. It is easy to see that if $\mathcal{T}_{n}\models r_{\tau}(\bar{a},b)$ and $\mathcal{T}^{\prime\prime}$ is a subtree of $\mathcal{T}_{n}$ which is rooted in $b$ and isomorphic to $\mathcal{T}^{\prime}$, then the elements of $T^{\prime\prime}\setminus\{b\}$ can be ordered as $\bar{c}$ so that $\mathcal{T}_{n}\models p_{\tau}(\bar{a},b,\bar{c})$. Also, note that if $\mathcal{T}^{*}$ is a different (from $\mathcal{T}^{\prime\prime}$) subtree of $\mathcal{T}_{n}$ that is rooted in $b$ and isomorphic to $\mathcal{T}^{\prime}$, then $T^{\prime\prime}\setminus\{b\}\neq T^{*}\setminus\{b\}$. Therefore, $|p_{\tau}(\bar{a},b,\mathcal{T}_{n})|$ is at least as large as the number of subtrees of $\mathcal{T}_{n}$ that are rooted in $b$ and isomorphic to $\mathcal{T}^{\prime}$. Recall that $\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})$ is the number of subtrees of $\mathcal{T}_{n}$ that are rooted in $b$ and isomorphic to $\mathcal{T}^{\prime}$. Now we have $|p_{\tau}(\bar{a},b,\mathcal{T}_{n})|\geq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})$ if $p_{\tau}(\bar{a},b,\mathcal{T}_{n})\neq\emptyset$. Recall that, by Remark 6.4, we have $\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})\geq g_{1}(n)$ whenever $n$ is large enough and $\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})>0$.
@@
We now assume that $p_{\tau}(\bar{a},\mathcal{T}_{n})\neq\emptyset$; otherwise there is nothing to prove. Then $r_{\tau}(\bar{a},\mathcal{T}_{n})\neq\emptyset$ and for every $b\in r_{\tau}(\bar{a},\mathcal{A})$ we have $|p_{\tau}(\bar{a},b,\mathcal{T}_{n})|\geq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})\geq g_{1}(n)$ if $n$ is large enough. We will now define a partition $P_{1},\ldots,P_{\mu}$ of $r_{\tau}(\bar{a},\mathcal{T}_{n})$. Recall the functions $g_{1},g_{2},g_{3}:\mathbb{N}\to\mathbb{R}^{+}$ from Assumption 6.3.
@@
    Let $\lambda\geq 1$ be the minimal integer such that there is $b\in r_{\tau}(\bar{a},\mathcal{A})$ such that $g_{1}(n)+(\lambda-1)g_{2}(n)\leq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})<g_{1}(n)+\lambda g_{2}(n)$.
@@
      If $\lambda=1$ then let $P_{1}$ contain all $b\in r_{\tau}(\bar{a},\mathcal{T}_{n})$ such that $m_{1,1}\leq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})<m_{1,2}$ where $m_{1,1}:=g_{1}(n)$ and $m_{1,2}:=m_{1,1}+(\lambda+3)g_{2}(n)$.
@@
      If $\lambda>1$ then let $P_{1}$ contain all $b\in r_{\tau}(\bar{a},\mathcal{T}_{n})$ such that $m_{1,1}\leq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})<m_{1,2}$ where $m_{1,1}:=g_{1}(n)+(\lambda-1)g_{2}(n)$ and $m_{1,2}:=m_{1,1}+(\lambda+3)g_{2}(n)$.
@@
    Now suppose that the parts $P_{1},\ldots,P_{l}$ have been defined such that $|P_{i}|\geq g_{3}(n)$ for all $i$ and that numbers $m_{i,j}$ (with $j\in\{1,2\}$) have been defined such that $P_{i}$ contains all $b\in r_{\tau}(\bar{a},\mathcal{A})$ such that $m_{i,1}\leq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})<m_{i,2}$, and $m_{i+1,1}=m_{i,2}$ if $i<l$.
    Let $\lambda\geq 1$ be the minimal integer (if it exists) such that there is $b\in r_{\tau}(\bar{a},\mathcal{A})$ such that $m_{l,2}+\lambda g_{2}(n)\leq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})<m_{l,2}+(\lambda+1)g_{2}(n)$.
@@
      If $\lambda=1$ then let $P_{l+1}$ contain all $b\in r_{\tau}(\bar{a},\mathcal{A})$ such that $m_{l+1,1}\leq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})<m_{l+1,2}$ where $m_{l+1,1}:=m_{l,2}$ and $m_{l+1,2}:=m_{l,2}+3g_{2}(n)$.
@@
        let $P_{l+1}$ contain all $b\in r_{\tau}(\bar{a},\mathcal{T}_{n})$ such that $m_{l+1,1}\leq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})<m_{l+1,2}$ where $m_{l+1,1}:=m_{l,2}+(\lambda-1)g_{2}(n)$ and $m_{l+1,2}:=m_{l,2}+(\lambda+2)g_{2}(n)$, and
@@
        redefine $P_{l}$ so that it contains all $b\in r_{\tau}(\bar{a},\mathcal{T}_{n})$ such that $m_{l,1}\leq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})<m_{l,2}+g_{2}(n)$ and then redefine $m_{l,2}$ by $m_{l,2}:=m_{l,2}+g_{2}(n)$.
@@
Observe that when a part $P_{i}$ was first defined then, for some real number $m_{i}$, all $b\in P_{i}$ satisfy $m_{i}\leq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})<m_{i}+3g_{2}(n)$. Also, each part $P_{i}$ was redefined at most once and in such a way that (with the same $m_{i}$) all $b\in P_{i}$ satisfy $m_{i}\leq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})<m_{i}+4g_{2}(n)$. By the construction we have $m_{i}\geq g_{1}(n)$ for all $i$. It follows that
@@
for all $b\in P_{i}$, $m_{i}\leq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})\leq m_{i}+4g_{2}(n)$.
@@
Recall that (by Assumption 6.3) every element of $\mathcal{T}_{n}$ has at most $g_{4}(n)$ children, where $g_{4}$ is a polynomial. Since all members of $r_{\tau}(\bar{a},\mathcal{T}_{n})$ are children of $a_{1}$ it follows that $|r_{\tau}(\bar{a},\mathcal{T}_{n})|\leq g_{4}(n)$. As $P_{1},\ldots,P_{\mu}$ is a partition of $r_{\tau}(\bar{a},\mathcal{T}_{n})$ it follows that $\mu\leq g_{4}(n)$. Let
@@
Proof. Let $\mathcal{A}\in\mathbf{X}_{n}^{\varepsilon}$, so $\mathcal{A}\in\mathbf{X}_{n}^{\varepsilon,i}$ for all $i=1,\ldots,\mu$, and let $\bar{a}\in(T_{n})^{|\bar{x}|}$. Suppose that $r_{\tau}(\bar{a},\mathcal{A})\neq\emptyset$, since otherwise the inequalities are trivial. Fix any $i\in\{1,\ldots,\mu\}$. By (7.9) there is a number $m_{i}\geq g_{1}(n)$ such that if $b\in P_{i}$ then $m_{i}\leq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})<m_{i}+4g_{2}(n)$ and hence $m_{i}\leq|p_{\tau}(\bar{a},b,\mathcal{A})|<m_{i}+4g_{2}(n)$.
@@
Let $p(\bar{x},\bar{y})$ and $r(\bar{x},\bar{y})$ be (possibly incomplete) closure types over $\sigma$, where $r$ is $\bar{y}$-positive with respect to $\mathbb{G}$, and let $q(\bar{x})$ be a complete closure type over $\sigma$, Then $(p,r,q)$ is balanced with respect to $\mathbb{G}$.
@@
Let $p(\bar{x},\bar{y})$, $r(\bar{x},\bar{y})$, and $q(\bar{x})$ be as in Proposition 7.24 and suppose that $p\wedge r\wedge q$ is cofinally satisfiable. By analysing the proofs of Section 7.4 (this section) we see that if $(p,r,q)$ is 00-balanced, then $(p,r)$ converges to 00. If $(p,r)$ is also eventually constant, then it must be eventually constant with value 00. This implies that for all sufficiently large $n$, all $\bar{a}\in(T_{n})^{|\bar{x}|}$, and all $\mathcal{A}\in\mathbf{W}_{n}$, we have $p(\bar{a},\mathcal{A})\cap r(\bar{a},\mathcal{A})=\emptyset$.